OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) <-> 2x − 6

myininaya (myininaya):

@darnelle Do you know how to distribute? a(b+c)=ab+ac Try that :)

OpenStudy (anonymous):

not really

myininaya (myininaya):

Well I showed you right there a(b+c) = ab+ac This is the distribute property Apply that to 2(x-3) and tell me what you get